OpenStudy (anonymous):

write an equation of the line that passes through the given points. (0,5),(1.5,1)

OpenStudy (anonymous):

Find the slope Slope=(y2−y1)/(x2−x1) and then use the slope and one of the given points and use the point-slope formula to get the equation of the line. y−y1=m(x−x1)

OpenStudy (anonymous):

IS THAT THE ANSWER I WANT THE ANSWER?

OpenStudy (anonymous):

That's an explanation of how you can find the answer. Let's work on it.

OpenStudy (anonymous):

ok let work on it then

OpenStudy (anonymous):

OK. Can you compute the slope?

OpenStudy (anonymous):

what that?

OpenStudy (anonymous):

Slope=(y2−y1)/(x2−x1)

OpenStudy (anonymous):

kk gotcha